You own a portfolio that has $1950 invested in Stock A and $3750 invested in Stock B. If the expected return on these stocks are 9 percent and 18 percent, respectively, what is the expected return on the portfolio?

a. 15.67%

b. 14.92%

c. 12.08

d. 15.22%

e. 13.50%
Answer #1

Total value = 1950 + 3750 = 5,700

Expected return on portfolio = (1950 / 5,700)*0.09 + (3750 / 5,700)*0.18

Expected return on portfolio = 0.030789 + 0.118421

Expected return on portfolio = 0.1492 or 14.92%